# Counting points on character varieties

### v1

@@ 5.4. Computing $\beta_{L,H}$
5.4.1.
We are reduced to computing the Möbius function and the functions $\beta_{L,H}^{\supseteq}$. The latter can be computed as follows. Consider the canonical quotient map to the group of coinvariants
@@
$\beta_{L,H}^{\supseteq}=\begin{cases}|\mathfrak{t}^{F}/\mathfrak{t}^{F}\cap[\mathfrak{l}^{F},\mathfrak{l}^{F}]|&\textrm{if $f_{\mathfrak{l}}(H)=0$; i.e., $H\in[\mathfrak{l}^{F},\mathfrak{l}^{F}]$;}\\ 0&\textrm{otherwise}.\end{cases}$
5.4.2.
@@
This is the additive analogue of (5). Observe that $\beta_{L,H}$ is independent of $H$ (provided it is generic).
@@ 5.5. Conclusion of the point count for once-punctured case
5.5.1.
It is clear from this description that $|{Y}^{F}|$ is a rational function in $q$. If we replace the Frobenius $F$ by $F^{m}$ for some positive integer $m$, then $|{Y}^{F^{m}}|$ is given by the same rational function with $q$ replaced by $q^{m}$. It follows that this rational function is actually a polynomial. Thus, ${Y}$ is polynomial count, in the sense of [LetellierVillegas], and our expression for $|{Y}^{F}|$ is the $E$-polynomial of $Y$. The fact that it is monic is left as an exercise to the reader; see [Giannini24] for details.
@@ 5.6. Multi-punctured case
5.6.1.
For the benefit of the reader, we recall what the notation means. Here, $Z$ is the centre of $G$, $\mathcal{T}(\mathfrak{g})$ is the set of $\mathfrak{g}$-types (i.e., $W$-orbits of pairs $(L,\mathcal{N})$, where $L$ is a standard Levi subgroup of $G$ and $\mathcal{N}$ a nilpotent $L^{F}$-orbit in $\mathfrak{l}^{F}$), and $d(\tau)=\dim(L)-\dim(L.n)$ where $n\in\mathcal{N}$, $Q_{T}^{L}$ is the Green function evaluated at the unipotent class $\varpi(\mathcal{N})\subseteq L^{F}$ associated to $\mathcal{N}$, $|[L]|$ is the size of the $W$-orbit of $L$, and $\mu(L,G)$ is Möbius function associated to the poset of standard Levi subgroups of $G$. The same argument as above implies that $Y$ is polynomial count with the monic counting polynomial given by $|Y^{F}|$; see [Giannini24] for details.
@@ 6. Examples
In this section, we give examples of Euler characteristics and counting polynomials of the character variety $X$ and its additive analogue $Y$. We also give tables containing $G$-types, $\mathfrak{g}$-types, and the associated data required to compute counting polynomials. In the case $(g,n)=(0,3)$, we also give counting polynomials when $G$ is a group of high rank such as $F_{4}$ and $E_{6}$ (other choices for $(g,n)$ are computable using the code found at [Bailey]).
Throughout, $\Phi_{i}$ is the $i$th cyclotomic polynomial. In particular,
@@
6.1. Euler characteristic of ${X}$
In this subsection, we give examples of Euler characteristics of the character variety $X$.
6.1.1. $g=1$ and trivial centre
Using Theorem 2, it is straightforward to show
$\chi({X})=|W|^{n-1}\sum_{L}|W(L)|\,|\mathrm{Irr}(W(L))|\,\nu(L),$
where the sum is over all endoscopy groups of $G$ containing $T$.
6.1.2.
Below we give tables containing the data required for computing this Euler characteristic in the cases $G=\mathrm{SO}_{5}$ and $G=G_{2}$. Using these tables, we find
$\chi({X}_{\mathrm{SO}_{5}})=72\times 8^{n-1}\qquad\text{and}\qquad\chi({X}_{G_{2}})=96\times 12^{n-1}.$
@@ 6.1.1. $g=1$ and trivial centre
Figure 2. Required data for $\chi(X)$ when $G=\mathrm{SO}_{5}$ and $g=1$.
@@
Figure 3. Required data for $\chi(X)$ when $G=G_{2}$ and $g=1$.
6.1.3. $g=0$
Using Theorem 2, it is straightforward to show the Euler characteristic is
$\chi({X})=\frac{1}{(2r)!}\frac{d^{2r}}{dq^{2r}}\bigg{|}_{q=1}\xi(q),$
where $r:=\dim(T)-\dim(Z)$ is the semisimple rank of $G$ and
$\xi(q):=q^{|\Phi(G)^{+}|(n-2)}\sum_{L}\bigg{(}\frac{|W|}{|W(L)|}\bigg{)}^{n-1}\nu(L)\sum_{\rho}\dim(\rho)^{n}\bigg{(}\frac{P_{W(L)}(q)}{\dim(\tilde{\rho})}\bigg{)}^{n-2}.$
Above, the first sum is over all endoscopy groups of $G$ containing $T$, and the second sum is over all irreducible characters of $W(L)$. We give a few examples of Euler characteristics below. Note that the Euler characteristic of ${X}$ is 00 if $n=1,2$, which is consistent with the fact that ${X}$ is empty in these cases.
$G$ $\chi({X})$
$\mathrm{GL}_{2}$ $2^{n-4}(n-1)(n-2)$
$\mathrm{GL}_{3}$ $2^{n-5}3^{n-3}(n-1)(n-2)(9n^{2}-27n+16)$
$\mathrm{SO}_{5}$ $2^{3n-8}(n-1)(n-2)(11n^{2}-33n+19)$
$G_{2}$ $2^{2n-7}3^{n-3}(n-1)(n-2)(207n^{2}-621n+350)$
Figure 4. Euler characteristics of ${X}$ when $g=0$.
6.2. Euler characteristic of ${Y}$
We now consider the additive character variety of Theorem 5. Using the counting polynomial provided by the theorem, it is straightforward to show
$\chi({Y})=\frac{1}{r!}\frac{d^{r}}{dq^{r}}\bigg{|}_{q=1}\zeta(q),$
where $r:=\dim(T)-\dim(Z)$ is the semisimple rank of $G$ and
$\zeta(q):=\frac{|\mathfrak{g}(k)|^{g-1}}{q^{|\Phi(G)^{+}|}P_{W}(q)}\sum_{\tau\in\mathcal{T}(\mathfrak{g})}q^{gd(\tau)}H_{\tau}(q).$
6.2.1.
We give a few examples of Euler characteristics below. As before, the Euler characteristic of ${Y}$ is 0 if $(g,n)=(0,1)$ or $(0,2)$ since $Y$ is empty in these cases.
$G$ $\chi({Y})$
@@ 6.2. Euler characteristic of ${Y}$
Figure 5. Euler characteristics of ${Y}$.
6.3. The group $G=\mathrm{SO}_{5}$
@@
$[{A_{1}\times A_{1}},2^{1}\otimes 2^{1}]$ $q^{4}\Phi_{1}^{2}\Phi_{2}^{2}$ $2^{n}\Phi_{1}^{2}$ $q^{2}\Phi_{1}^{2}\Phi_{2}^{2}$ $1$ $1$ $4$ $1$ $4$ $2$
$[{A_{1}\times A_{1}},2^{1}\otimes 1^{2}]$ $q^{3}\Phi_{1}^{2}\Phi_{2}^{2}$ $2^{n}\Phi_{1}^{2}$ $q^{2}\Phi_{1}^{2}\Phi_{2}^{2}$ $q$ $1$ $4$ $1$ $4$ $2$
$[{A_{1}\times A_{1}},1^{2}\otimes 2^{1}]$ $q^{3}\Phi_{1}^{2}\Phi_{2}^{2}$ $2^{n}\Phi_{1}^{2}$ $q^{2}\Phi_{1}^{2}\Phi_{2}^{2}$ $q$ $1$ $4$ $1$ $4$ $2$
$[{A_{1}\times A_{1}},1^{2}\otimes 1^{2}]$ $q^{2}\Phi_{1}^{2}\Phi_{2}^{2}$ $2^{n}\Phi_{1}^{2}$ $q^{2}\Phi_{1}^{2}\Phi_{2}^{2}$ $q^{2}$ $1$ $4$ $1$ $4$ $2$
$[{A_{1}},2^{1}]$ $q^{4}\Phi_{1}^{2}\Phi_{2}$ $-2\cdot 4^{n}\Phi_{1}^{2}$ $q\Phi_{1}^{2}\Phi_{2}$ $1$ $1$ $2$ $2$ $2$ $-4$
$[{A_{1}},1^{2}]$ $q^{3}\Phi_{1}^{2}\Phi_{2}$ $-2\cdot 4^{n}\Phi_{1}^{2}$ $q\Phi_{1}^{2}\Phi_{2}$ $q$ $1$ $2$ $2$ $2$ $-4$
$[{A_{1}^{\prime}},2^{1}]$ $q^{4}\Phi_{1}^{2}\Phi_{2}$ $-4^{n}\Phi_{1}^{2}$ $q\Phi_{1}^{2}\Phi_{2}$ $1$ $1$ $2$ $2$ $1$ $-2$
$[{A_{1}^{\prime}},1^{2}]$ $q^{3}\Phi_{1}^{2}\Phi_{2}$ $-4^{n}\Phi_{1}^{2}$ $q\Phi_{1}^{2}\Phi_{2}$ $q$ $1$ $2$ $2$ $1$ $-2$
$[T,1^{1}]$ $q^{4}\Phi_{1}^{2}$ $8^{n}\Phi_{1}^{2}$ $\Phi_{1}^{2}$ $1$ $1$ $1$ $1$ $1$ $8$
Figure 6. The fourteen $\mathrm{SO}_{5}$-types.
$[\mathrm{SO}_{5},1^{5}]$ $q^{10}$ $q^{4n}\Phi_{2}^{2n}\Phi_{4}^{n}$ $q^{4}\Phi_{1}^{2}\Phi_{2}^{2}\Phi_{4}$ $1$ $\Phi_{2}^{2}\Phi_{4}$ $8$ $1$ $1$
$[\mathrm{SO}_{5},2^{2}1^{1}]$ $q^{6}$ $q^{4n}\Phi_{1}\Phi_{2}^{2n+1}\Phi_{4}$ $q^{4}\Phi_{1}^{2}\Phi_{2}^{2}\Phi_{4}$ $\Phi_{1}\Phi_{2}\Phi_{4}$ $\Phi_{2}^{2}$ $8$ $1$ $1$
$[\mathrm{SO}_{5},3^{1}1^{2}]$ $q^{4}$ $\frac{1}{2}q^{4n+1}\Phi_{1}\Phi_{2}^{2}\Phi_{4}(3q+1)^{n}$ $q^{4}\Phi_{1}^{2}\Phi_{2}^{2}\Phi_{4}$ $\frac{1}{2}q\Phi_{1}\Phi_{2}^{2}\Phi_{4}$ $13q+$ 1 $8$ $1$ $1$
$[\mathrm{SO}_{5},3^{1}1^{2}_{\star}]$ $q^{4}$ $\frac{1}{2}q^{4n+1}\Phi_{1}^{2}\Phi_{2}^{n+1}\Phi_{4}$ $q^{4}\Phi_{1}^{2}\Phi_{2}^{2}\Phi_{4}$ $\frac{1}{2}q\Phi_{1}^{2}\Phi_{2}\Phi_{4}$ $\Phi_{2}$ $8$ $1$ $1$
$[\mathrm{SO}_{5},5^{1}]$ $q^{2}$ $q^{4n+2}\Phi_{1}^{2}\Phi_{2}^{2}\Phi_{4}$ $q^{4}\Phi_{1}^{2}\Phi_{2}^{2}\Phi_{4}$ $q^{2}\Phi_{1}^{2}\Phi_{2}^{2}\Phi_{4}$ $1$ $8$ $1$ $1$
$[A_{1},1^{2}]$ $q^{4}$ $-2\cdot 4^{n-1}q^{4n+3}\Phi_{2}^{n+1}\Phi_{4}$ $q\Phi_{1}^{2}\Phi_{2}$ $1$ $\Phi_{2}$ $2$ $2$ $-1$
$[A_{1},2^{1}]$ $q^{2}$ $-2\cdot 4^{n-1}q^{4n+3}\Phi_{1}\Phi_{2}^{2}\Phi_{4}$ $q\Phi_{1}^{2}\Phi_{2}$ $\Phi_{1}\Phi_{2}$ $1$ $2$ $2$ $-1$
$[A_{1}^{\prime},1^{2}]$ $q^{4}$ $-2\cdot 4^{n-1}q^{4n+3}\Phi_{2}^{n+1}\Phi_{4}$ $q\Phi_{1}^{2}\Phi_{2}$ $1$ $\Phi_{2}$ $2$ $2$ $-1$
$[A_{1}^{\prime},2^{1}]$ $q^{2}$ $-2\cdot 4^{n-1}q^{4n+3}\Phi_{1}\Phi_{2}^{2}\Phi_{4}$ $q\Phi_{1}^{2}\Phi_{2}$ $\Phi_{1}\Phi_{2}$ $1$ $2$ $2$ $-1$
$[T,0]$ $q^{2}$ $3\cdot 8^{n-1}q^{4n+4}\Phi_{2}^{2}\Phi_{4}$ $\Phi_{1}^{2}$ $1$ $1$ $1$ $1$ $3$
Figure 7. The ten $\mathfrak{so}_{5}$-types.
$(g,n)$ $|{X}(\mathbb{F}_{q})|$
@@
Figure 8. Examples of the polynomial $|{X}(\mathbb{F}_{q})|$ when $G=\mathrm{SO}_{5}$.
$(g,n)$ $|{Y}(\mathbb{F}_{q})|$

### v2

@@ 6.4. Computing $\beta_{L,H}$
6.4.1.
We are reduced to computing the Möbius function and the functions $\beta_{L,H}^{\supseteq}$. The latter can be computed as follows. Consider the canonical quotient map from $\mathfrak{t}^{F}$ to the group of $W(L)$-coinvariants
@@
$\beta_{L,H}^{\supseteq}=\begin{cases}|\mathfrak{t}^{F}/\mathfrak{t}^{F}\cap[\mathfrak{l}^{F},\mathfrak{l}^{F}]|&\textrm{if $H\in[\mathfrak{l}^{F},\mathfrak{l}^{F}]$;}\\ 0&\textrm{otherwise}.\end{cases}$
6.4.2.
@@
This is the additive analogue of Proposition 32. Observe that $\beta_{L,H}$ is independent of $H$ (assuming $H$ is generic).
@@ 6.5. Conclusion of the point count for once-punctured case
6.5.1.
It is shown in [GoodwinRohrle, Theorem 3.10] that (under the assumption that $Z$ is connected and $p$ is good for $G$) the Green function $Q_{T}^{G}(u)$ is a polynomial in $q$, unless $G$ has a factor of type $E_{8}$, in which case a quadratic base change may be necessary. We can rephrase this by saying that for every type $\tau=[L,\mathcal{N}]$, there exists a polynomial $f_{\tau}\in\mathbb{Z}[t]$ such that the size of the Springer fibre $Q_{T}^{L}(\varpi(\mathcal{N}))$ over the finite field $\mathbb{F}_{q^{2r}}$ equals $f_{\tau}(q^{2r})$ for all positive integers $r$.
6.5.2.
Let $s_{\tau}\in\mathbb{Z}[t]$ denote the counting polynomial for $C_{L}(n)$ where $n\in\mathcal{N}$. Let
$\gamma_{G}:=\dim(Z)+(g-1)\dim(\mathfrak{g})+|\Phi^{+}|.$
Define the rational function $f_{Y}$ by
$f_{Y}(t):=f_{Z}(t)t^{\gamma_{G}}\sum_{\tau\in\mathcal{T}(\mathfrak{g})}t^{gd(\tau)}\frac{f_{\tau}(t)}{s_{\tau}(t)}\mu(L,G)|[L]|\in\mathbb{Q}(t),$
where $f_{Z}(t)=(t-1)^{\dim(Z)}$ is the counting polynomial for $Z$ as noted in §4.5.1. The above discussions imply
$|Y(\mathbb{F}_{q^{2r}})|=f_{Y}(q^{2r}),\qquad\forall r\in\mathbb{Z}_{>0}.$
6.5.3.
By [GoodwinRohrle, Lemma 2.11], $f_{Y}\in\mathbb{Q}[t]$. Thus, $Y\times_{\mathbb{F}_{q}}\mathbb{F}_{q^{2}}$ is polynomial count with counting polynomial $f_{Y}$. This concludes the proof of Theorem 6 in the single punctured case.
@@ 6.6. Multi-punctured case
6.6.1.
Now, let
$\gamma_{G}:=\dim(Z)+(g-1)\dim(\mathfrak{g})+n|\Phi^{+}|.$
Define the rational function $f_{Y}$ by
$f_{Y}(t):=f_{Z}(t)t^{\gamma_{G}}\sum_{\tau\in\mathcal{T}(\mathfrak{g})}t^{gd(\tau)}\bigg{(}\frac{|W|}{|W(L)|}\bigg{)}^{n-1}\frac{(f_{\tau}(t))^{n}}{s_{\tau}(t)}\mu(L,G)|[L]|\in\mathbb{Q}(t).$
Then
$|Y(\mathbb{F}_{q^{2r}})|=f_{Y}(q^{2r}),\qquad\forall r\in\mathbb{Z}_{>0}.$
6.6.2.
The same argument as the previous subsection implies that $Y\times_{\mathbb{F}_{q}}\mathbb{F}_{q^{2}}$ is polynomial count with counting polynomial $f_{Y}$. This concludes the proof of Theorem 6.
6.7. Connectedness
Our aim now is to show that $\overline{Y}$ is connected. We know that $\overline{Y}$. Thus, it suffices to show that the $E$-polynomial $f_{Y}$ is monic. The $\mathfrak{g}$-type contributing to the leading term of $f_{Y}$ is $\tau_{1}=[G,0]$. The coefficient of this term is determined by the Green polynomial $Q_{T}^{G}(1)$ because
$H_{\tau_{1}}=q^{n|\Phi^{+}|+\dim(Z)}Q_{T}^{G}(\varpi(0)).$
The term $Q_{T}^{G}(\varpi(0))$ is the Poincaré polynomial of $G$ and is known to be monic. Thus, the polynomial $c(q):=\sum_{\tau\in\mathcal{T}(\mathfrak{g})}q^{gd(\tau)}H_{\tau}(q)$ is monic. Now,
$|Y(\mathbb{F}_{q^{2r}})|=\frac{|Z(\mathbb{F}_{q^{2r}})||\mathfrak{g}(\mathbb{F}_{q^{2r}})|}{|G(\mathbb{F}_{q^{2r}})|}c(q)$
is the quotient of two monic polynomials, therefore it is also monic.∎
6.8. Euler characteristic
It follows from the explicit description of $f_{Y}$ that
$f_{Y}(1)=\frac{1}{r!}\frac{d^{r}}{dq^{r}}\bigg{|}_{q=1}\zeta(q),$
where $r:=\dim(T)-\dim(Z)$ is the semisimple rank of $G$ and
$\zeta(q):=\frac{|\mathfrak{g}(k)|^{g-1}}{q^{|\Phi^{+}|}P_{W}(q)}\sum_{\tau\in\mathcal{T}(\mathfrak{g})}q^{gd(\tau)}H_{\tau}(q),$
where $P_{W}(q)$ is the Poincaré polynomial of $W$. We use this formula to give examples of $\chi(\overline{Y})$ in the next section.
@@ 7. Examples
In this section, we give examples of Euler characteristics and counting polynomials of generic character varieties. We also give tables containing $G$-types, $\mathfrak{g}$-types, and the associated data required to compute the counting polynomials. Some expressions are obtained with the help of the computer; see [Bailey] for the computer code which is based on [GHLMP96].
Throughout, $\Phi_{i}$ denotes the $i$th cyclotomic polynomial; thus, for instance
@@
7.1. Euler characteristics for $\overline{X}$ in genus 00
$G$ $\chi(\overline{{X}})$
$\mathrm{GL}_{2}$ $2^{n-4}(n-1)(n-2)$
$\mathrm{GL}_{3}$ $2^{n-5}3^{n-3}(n-1)(n-2)(9n^{2}-27n+16)$
$\mathrm{SO}_{5}$ $2^{3n-8}(n-1)(n-2)(11n^{2}-33n+19)$
$G_{2}$ $2^{2n-7}3^{n-3}(n-1)(n-2)(207n^{2}-621n+350)$
Note that $\chi(\overline{X})=0$ when $n=1$ or $2$, in agreement with the fact that the (generic) character variety is empty in these cases.
7.2. Euler characteristics of $\overline{X}$ when $(G,g)=(\mathrm{SO}_{5},1)$
@@
The above table displays the required data for computing $\chi(\overline{X})$ when $G=\mathrm{SO}_{5}$ and $g=1$. Using this table, we find ${\chi(\overline{{X}})=72\times 8^{n-1}}$.
7.3. Euler characteristics of $\overline{X}$ when $(G,g)=(G_{2},1)$
@@
The above table displays the required data for computing $\chi(\overline{X})$ when $G=G_{2}$ and $g=1$. Using this table, we find $\chi(\overline{{X}})=96\times 12^{n-1}$.
7.4. Euler characteristic of $\overline{{Y}}$
$G$ $\chi(\overline{Y})$
@@ 7.4. Euler characteristic of $\overline{{Y}}$
7.5. Types for $G=\mathrm{SO}_{5}$ and expressions for the counting polynomial $f_{X}$
@@
$[{A_{1}\times A_{1}},2^{1}\otimes 2^{1}]$ $q^{4}\Phi_{1}^{2}\Phi_{2}^{2}$ $2^{n}$ $q^{2}\Phi_{1}^{2}\Phi_{2}^{2}$ $1$ $1$ $4$ $1$ $4$ $2$
$[{A_{1}\times A_{1}},2^{1}\otimes 1^{2}]$ $q^{3}\Phi_{1}^{2}\Phi_{2}^{2}$ $2^{n}$ $q^{2}\Phi_{1}^{2}\Phi_{2}^{2}$ $q$ $1$ $4$ $1$ $4$ $2$
$[{A_{1}\times A_{1}},1^{2}\otimes 2^{1}]$ $q^{3}\Phi_{1}^{2}\Phi_{2}^{2}$ $2^{n}$ $q^{2}\Phi_{1}^{2}\Phi_{2}^{2}$ $q$ $1$ $4$ $1$ $4$ $2$
$[{A_{1}\times A_{1}},1^{2}\otimes 1^{2}]$ $q^{2}\Phi_{1}^{2}\Phi_{2}^{2}$ $2^{n}$ $q^{2}\Phi_{1}^{2}\Phi_{2}^{2}$ $q^{2}$ $1$ $4$ $1$ $4$ $2$
$[{A_{1}},2^{1}]$ $q^{4}\Phi_{1}^{2}\Phi_{2}$ $-2\cdot 4^{n}$ $q\Phi_{1}^{2}\Phi_{2}$ $1$ $1$ $2$ $2$ $2$ $-4$
$[{A_{1}},1^{2}]$ $q^{3}\Phi_{1}^{2}\Phi_{2}$ $-2\cdot 4^{n}$ $q\Phi_{1}^{2}\Phi_{2}$ $q$ $1$ $2$ $2$ $2$ $-4$
$[{A_{1}^{\prime}},2^{1}]$ $q^{4}\Phi_{1}^{2}\Phi_{2}$ $-4^{n}$ $q\Phi_{1}^{2}\Phi_{2}$ $1$ $1$ $2$ $2$ $1$ $-2$
$[{A_{1}^{\prime}},1^{2}]$ $q^{3}\Phi_{1}^{2}\Phi_{2}$ $-4^{n}$ $q\Phi_{1}^{2}\Phi_{2}$ $q$ $1$ $2$ $2$ $1$ $-2$
$[T,1^{1}]$ $q^{4}\Phi_{1}^{2}$ $8^{n}$ $\Phi_{1}^{2}$ $1$ $1$ $1$ $1$ $1$ $8$
The above table displays $\mathrm{SO}_{5}$-types and the data require for counting points on character varieties. The point counts for low values of $g$ and $n$ are given explicitly below. Since the polynomials are palindromic, we only give “half” of them.
$(g,n)$ $f_{X}(q)$
@@
7.6. Types for $\mathfrak{so}_{5}$ and expressions for the counting polynomial $f_{Y}$
$\tau=[L,\mathcal{N}]$ ${d(\tau)}$ $H_{\tau}(q)$ $|L^{F}|$ $|\mathcal{N}|$ $Q_{T}^{L}(\varpi(\mathcal{N}))$ $|W(L)|$ $|[L]|$ $\mu(L,G)$
$[\mathrm{SO}_{5},1^{5}]$ 1010{10}10 $q^{4n}\Phi_{2}^{2n}\Phi_{4}^{n}$ $q^{4}\Phi_{1}^{2}\Phi_{2}^{2}\Phi_{4}$ $1$ $\Phi_{2}^{2}\Phi_{4}$ $8$ $1$ $1$
$[\mathrm{SO}_{5},2^{2}1^{1}]$ 66{6}6 $q^{4n}\Phi_{1}\Phi_{2}^{2n+1}\Phi_{4}$ $q^{4}\Phi_{1}^{2}\Phi_{2}^{2}\Phi_{4}$ $\Phi_{1}\Phi_{2}\Phi_{4}$ $\Phi_{2}^{2}$ $8$ $1$ $1$
$[\mathrm{SO}_{5},3^{1}1^{2}]$ 44{4}4 $\frac{1}{2}q^{4n+1}\Phi_{1}\Phi_{2}^{2}\Phi_{4}(3q+1)^{n}$ $q^{4}\Phi_{1}^{2}\Phi_{2}^{2}\Phi_{4}$ $\frac{1}{2}q\Phi_{1}\Phi_{2}^{2}\Phi_{4}$ $13q+$ 1 $8$ $1$ $1$
$[\mathrm{SO}_{5},3^{1}1^{2}_{\star}]$ 44{4}4 $\frac{1}{2}q^{4n+1}\Phi_{1}^{2}\Phi_{2}^{n+1}\Phi_{4}$ $q^{4}\Phi_{1}^{2}\Phi_{2}^{2}\Phi_{4}$ $\frac{1}{2}q\Phi_{1}^{2}\Phi_{2}\Phi_{4}$ $\Phi_{2}$ $8$ $1$ $1$
$[\mathrm{SO}_{5},5^{1}]$ 22{2}2 $q^{4n+2}\Phi_{1}^{2}\Phi_{2}^{2}\Phi_{4}$ $q^{4}\Phi_{1}^{2}\Phi_{2}^{2}\Phi_{4}$ $q^{2}\Phi_{1}^{2}\Phi_{2}^{2}\Phi_{4}$ $1$ $8$ $1$ $1$
$[A_{1},1^{2}]$ 44{4}4 $-2\cdot 4^{n-1}q^{4n+3}\Phi_{2}^{n+1}\Phi_{4}$ $q\Phi_{1}^{2}\Phi_{2}$ $1$ $\Phi_{2}$ $2$ $2$ $-1$
$[A_{1},2^{1}]$ 22{2}2 $-2\cdot 4^{n-1}q^{4n+3}\Phi_{1}\Phi_{2}^{2}\Phi_{4}$ $q\Phi_{1}^{2}\Phi_{2}$ $\Phi_{1}\Phi_{2}$ $1$ $2$ $2$ $-1$
$[A_{1}^{\prime},1^{2}]$ 44{4}4 $-2\cdot 4^{n-1}q^{4n+3}\Phi_{2}^{n+1}\Phi_{4}$ $q\Phi_{1}^{2}\Phi_{2}$ $1$ $\Phi_{2}$ $2$ $2$ $-1$
$[A_{1}^{\prime},2^{1}]$ 22{2}2 $-2\cdot 4^{n-1}q^{4n+3}\Phi_{1}\Phi_{2}^{2}\Phi_{4}$ $q\Phi_{1}^{2}\Phi_{2}$ $\Phi_{1}\Phi_{2}$ $1$ $2$ $2$ $-1$
$[T,0]$ 22{2}2 $3\cdot 8^{n-1}q^{4n+4}\Phi_{2}^{2}\Phi_{4}$ $\Phi_{1}^{2}$ $1$ $1$ $1$ $1$ $3$
The above table displays $\mathfrak{so}_{5}$-types and the data require for counting points on the additive character varieties. The point counts for low values of $g$ and $n$ are given explicitly below.
$(g,n)$ $f_{Y}(q)$